# A generalized Apollonius problem

### v1

@@ Abstract.
We present a self-contained solution of a generalized Apollonius problem of constructing a circle that intersects each of the three pair-wise non-intersecting circles at one of the three given angles.
@@ 1. Introduction
In the $3^{\text{rd}}$ century BC, Apollonius of Perga posted and solved the famous problem of constructing – with the help of a straightadge and a compass only – a circle tangent to three given circles - see, for example, [4, 2, 3] for a general statement.
@@
The significance of this problem can hardly be overstated - no wonder the title of the Presidential address delivered in 1967 by the great geometer of the $20^{\text{th}}$ century H.S.M. Coxeter to the Canadian Mathematical Congress at the joint meeting with the Mathematical Association of America was ”The Problem of Apollonius.” At the beginning of his presentation Coxeter stated, ”As Sir Thomas Heath remarks [6, p. 182], this problem ’has exercised the ingenuity of many distinguished geometers, including Vieta and Newton.’” (The results presented in this address were later published in [2].)
As a result, there are many ingenious approaches to the Apollonius problem. Unfortunately, Apollonius’ solution of his problem has not survived. However, François Viète found a solution by considering limiting cases, where any of the three given circles can be reduced to a point or blown up into a line, which one may believe to be the Apollonius’ approach.
The method of inversion yields an efficient way of solving Apollonius problem, where two of the three given circles are transformed either into two parallel lines or into two concentric circles. This approach, in part, utilizes the fact that inversion preserves angles between two intersecting circles. We use this fact, together with the availability of an inversion that transforms two non-intersecting circles into concentric ones, to solve the problem of constructing – using a straightedge and a compass only – a cirle that intersects each of the three given circles at one of the three given angles from the interval $[0,90^{0}]$. The Apollonius problem is a particular case of this problem where the three given angles are zeros.
@@
To solve this problem, we first assume that such a circle exists and perform an inversion of the four circles that transforms two of the three given circles into concentric ones. Then we find a formula for a possible radius of the image of the circle in question and formulas for possible distances from its center to the centers of the images of the given circles. Finally, using these three formulas, we construct a circle $\mathcal{C}^{\prime}$ that does intersect each of the images of the three given ones at one of the three given angles. As the square of an inversion is the identity map, the inversion of $\mathcal{C}^{\prime}$ delivers a circle that intersects each of the three given circles at one of the three given angles.
@@ 2.2. Constructibility
Let us define what it means for an interval, or an angle, to be constructible.
Definition 2.1.
Fix a unit interval, and let $a_{1},a_{2},\dots,a_{n}$ be intervals. Denote their lengths by the same letters. An interval of length
(1) $a=f(a_{1},a_{2},\dots,a_{n})$
is called constructible if it can be drawn with the help of straightedge and a compass. In this case, the number $a$ is also called constructible. An angle $\alpha$ is called constructible if $\cos\alpha$ (equivalently, $\sin\alpha$) is a constructible number. If an interval (angle) can be constructed avoiding the use of the unit interval, we will call it properly constructible or p-consructible.
We presume that the reader is familiar with the following straightedge and compass p-constructions:
@@
    Drawing the line through a given point and perpendicular to a given line.
@@
    Drawing the line through a given point and parallel to a given line.
Example 2.1.
P-construct the circle that passes through three non-collinear points $A$, $B$, and $C$.
Solution: Utilizing constructions (a) and (b), draw perpendicular bisectors of the intervals $[AB]$ and $[BC]$. The point of intersection of these bisectors is the center of the circle though $A$, $B$, and $C$.
To outline the scope of constructible intervals it is sufficient to consider a few examples where we use the constructions (a) - (c) without explicit references.
Example 2.2.
Given intervals $a$ and $b$, p-construct an interval of length $x=\sqrt{ab}$.
Solution: Draw a segment $[AB]$ and a point $C\in[AB]$ such that $|AC|=a$ and $|CB|=b$ as in Figure 2. Then draw a semi-circle with $[AB]$ as its diameter. Next, construct the line through $C$ perpendicular to the line $AB$. If $X$ is the point of intersection of this line with the semi-circle, then $\angle AXB=90^{0}$, implying that the triangles $ACX$ and $XCB$ are similar. Therefore, denoting $|CX|=x$, we have
$\frac{|CX|}{|AC|}=\frac{|CB|}{|CX|}\ \ \Longleftrightarrow\ \ \frac{x}{a}=\frac{b}{x}\ \ \Longleftrightarrow\ \ x=\sqrt{ab},$
thus, the interval $[CX]$ has the desired length.
Figure 2
Remark 2.1.
In particular, if $b=1$, then $x=\sqrt{a}$, that is, square roots of constructible intervals are also constructible.
Example 2.3.
Given intervals $a$, $b$, and $c$, p-construct an interval of length $x=\displaystyle\frac{ab}{c}$.
@@
In particular, when $c=1$ ($b=1$), we have $x=ab$ (respectively, $x=\displaystyle\frac{a}{c}$ ), that is, the products or ratios of constructible intervals are constructible.
Example 2.4.
Given intervals $a$, $b$, $c$, $d$, and an angle $\alpha\in(0,\pi)$, construct an interval of length
$e=\frac{\sqrt{a^{2}+bc\sin\alpha}}{\sqrt[4]{d}}.$
Solution: In view of Remark 2.2, one can construct an interval of length $a^{2}$ and an interval of length $bc$, hence an interval of length $bc\sin\alpha$. Thus, an interval of length $x=a^{2}+bc\sin\alpha$ can be constructed. Next, in view of Remark 2.1, one can construct an interval of length $\sqrt{x}=\sqrt{a^{2}+bc\sin\alpha}$. Further, as $4]{d}=\sqrt{\sqrt{d}}$, Remark 2.1 yields an interval of length $y=\sqrt[4]{d}$. Finally applying Remark 2.2, we end up with an interval of length $\frac{\sqrt{x}}{y}=e$.
2.3. Criterion of Constructibility
We can now state that if the function $f$ in (1) can be written as involving only the four arithmetic operations and operations of finding square roots, then the interval $a$ is constructible. In fact, it can be proven, as it is done in [1, pp. 129-134], that the converse statement is also true.
3. Statement of the Problem
3.1. Apollonius Three Circle Problem
Referring to Figure 4, we can state the famous Apollonius problem as follows:
@@ 3.1. Apollonius Three Circle Problem
Given three circles $\mathcal{C}_{1}$, $\mathcal{C}_{2}$, $\mathcal{C}_{3}$, p-construct a circle $\mathcal{C}$ that is tangent to each of them, that is, intersects each at one point.
@@
Note that, generically, there are eight different solutions to Apollonius problem, as the circle in question may touch one, or two, or all three given circles from outside. Moreover, the circles $\mathcal{C}_{1}$, $\mathcal{C}_{2}$, $\mathcal{C}_{3}$ may not be positioned as conveniently as in Figure 4.
Referring to Figure 5, we propose the following, that is a generalization of the Apollonius problem:
@@
Given three circles  $\mathcal{C}_{1}$, $\mathcal{C}_{2}$, $\mathcal{C}_{3}$ with centers at $O_{1}$, $O_{2}$, $O_{3}$ and radii $r_{1}$, $r_{2}$, $r_{3}$, respectively, p-construct a circle $\mathcal{C}$ that intersects $\mathcal{C}_{1}$ at an angle $0^{0}\leq\alpha\leq 90^{0}$, $\mathcal{C}_{2}$ at an angle $0^{0}\leq\beta\leq 90^{0}$, and $\mathcal{C}_{3}$ at an angle $0^{0}\leq\gamma\leq 90^{0}$.
@@
4. Solution of Problem 3.2
An elegant solution of the Apollonius problem - in a particular case given by Figure 4 - is given via an application of inversion as follows. First, the radii of the three given circles are enlarged by the same interval - so that two of the resulting circles become tangent at, say, a point $O$ and then performing an inversion with the center $O$ resulting in these two circles turn into two parallel lines.
Our solution is based on the following facts:
  1. (A)
    An inversion preserves the angles between two circles - see [3, Theorem 5.51].
  2. (B)
    Given two non-intersecting circles $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$, there exists a circle $\mathcal{C}_{\mathcal{I}}$ such that $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ turn into concentric circles under the inversion with respect to $\mathcal{C}_{\mathcal{I}}$ see [3, Theorem 5.71]. Note that this property can be used to find more solutions to the Apollonius problem - see [5].
4.1. Inversion and its Properties
@@
Points $P\neq O$ and $P^{\prime}\neq O$ are called the inversion images of each other with respect to $\mathcal{C}_{\mathcal{I}}$ if $P$ and $P^{\prime}$ lie on the same ray aminating from $O$ and
@@
This transformation of the plane - minus the point $O$ is called the inversion with respect to $\mathcal{C}_{\mathcal{I}}$. If we denote this inversion by $\mathcal{I}$, then $\mathcal{I}(P)=P^{\prime}$, $\mathcal{I}(P^{\prime})=P$, hence $\mathcal{I}^{2}(P)=P$ for every $P\neq O$, and $\mathcal{I}(P)=P$ whenever $P\in\mathcal{C}_{\mathcal{I}}$.
Figure 6 gives a way of p-constructing the point $P^{\prime}$ if $P$ is given and vise versa: If $P$ is given, let $A$ be the intersection of $\mathcal{C}_{\mathcal{I}}$ with the circle for which $[OP]$ is a diameter. Then $PA$ is tangent to $\mathcal{C}_{\mathcal{I}}$, and if $AP^{\prime}\perp OP$, the triangles $AOP$ and $P^{\prime}OA$ are similar, so we have
@@
Next, it can be proven - see [3, Theorem 5.45] - that
@@
    If a circle $\mathcal{C}$ passes through $O$, then $\mathcal{I}(\mathcal{C})$ is a straigh line.
Let us now describe a construction that will need to be performed repeatedly in the sequel without explicit reference:
Example 4.1.
Given a circle $\mathcal{C}$ that does not pass through the center of an inversion circle $\mathcal{C}_{\mathcal{I}}$, p-construct $\mathcal{I}(\mathcal{C})$.
Solution: Pick three points $A,B,C$ on $\mathcal{C}$ and p-construct the points $\mathcal{I}(A)$, $\mathcal{I}(B)$, $\mathcal{I}(C)$ using the procedure given by Figure 6. Then p-construct the circle $\mathcal{I}(\mathcal{C})$ as passing through the points $\mathcal{I}(A)$, $\mathcal{I}(B)$, $\mathcal{I}(C)$ as in Example 2.1.
4.2. P-constructing a circle $\mathcal{C}_{\mathcal{I}}$ in (B)
Referring to [3, Ch. 5, § 7] and Figure 7, we can describe the process of p-constructing possible centers for the circle $\mathcal{C}_{\mathcal{I}}$ from the statement (B) above as follows.
@@ 4.2. P-constructing a circle $\mathcal{C}_{\mathcal{I}}$ in (B)
Let $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ be non-intersecting circles with centers at $O_{1}$ and $O_{2}$ and $radii$ $r_{1}$ and $r_{2}$, respectively. P-constructing a right triangle with the hypotenuse $[O_{1}O_{2}]$ and one of the sides of length $|r_{1}-r_{2}|$, we can draw the common outer tangents to the circles $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$. Denote by $A_{1}$, $B_{1}$ and $A_{2}$, $B_{2}$ the points of intersection of these tangents with circles $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$, respectively, and let $M_{1}$ and $M_{2}$ be the mid-points of segments $A_{1}B_{1}$ and $A_{2}B_{2}$, respectively. Next, denote by $C$ the point of intersection of the lines $O_{1}O_{2}$ and $M_{1}M_{2}$.
P-construct a point $D\in\mathcal{C}_{1}$, as we did in Figure 6 with point $A$, such that $CD$ is tangent to $\mathcal{C}_{1}$ and draw the circle of radius $|CD|$ with the center at $C$. Let this circle intersect the segment $[O_{1}O_{2}]$ at points $\mathcal{O}_{\mathcal{I}_{1}}$ and $\mathcal{O}_{\mathcal{I}_{2}}$ counting from $O_{1}$. According to [3], each of the points $\mathcal{O}_{\mathcal{I}_{1}}$ and $\mathcal{O}_{\mathcal{I}_{2}}$ can serve as the center of the inversion cirle in (B).
For the reader’s convenience, Figure 8 complements Figure 5 by depicting the circle $\mathcal{C}_{\mathcal{I}}$, in green, that we will use to ensure that $\mathcal{C}_{2}^{\prime}=\mathcal{I}(\mathcal{C}_{2})$ and $\mathcal{C}_{3}^{\prime}=\mathcal{I}(\mathcal{C}_{3})$ are concentric. Note that the result of this inversion, given in Figure 9, lies entirely inside $\mathcal{C}_{\mathcal{I}}$.

### v2

# A contemporary inversive geometry solution of the three-circle problem of Steiner

@@ Abstract.
We present a concise self-contained inversive geometry solution of the three-circle problem of Steiner of constructing a circle that intersects each of the three given circles at one of the three given angles.
@@ 1. Introduction
In the $3^{\text{rd}}$ century BC, Apollonius of Perga posted and solved the famous problem of constructing a circle tangent to three given circles, stated as Problem A in Subsection 5.1 of this article – see also, for example, [8, 5, 7]. (Figuratively speaking, a geometric object is called constructible if it can be drawn using given geometric objects with the help of a straightedge and a compass only). However, Apollonius’ solution the problem has not survived.
@@
The significance of this problem can hardly be overstated – no wonder the title of the Presidential address delivered in 1967 by the great geometer of the $20^{\text{th}}$ century H.S.M. Coxeter, at the joint meeting with the Mathematical Association of America, to the Canadian Mathematical Congress was “The Problem of Apollonius” At the beginning of his presentation Coxeter stated, “As Sir Thomas Heath remarks [10, p. 182], this problem ‘has exercised the ingenuity of many distinguished geometers, including Vieta and Newton.’” (The results presented in this address were later published in [5].) Of note, Vieta found a solution of the problem of Apollonius by considering limiting cases, where any of the three given circles can be reduced to a point or blown up into a line, which may have been the Apollonius’ original approach.
This article is devoted to solving a generalized Apollonius problem of constructing a circle that intersects each of the three given circles at one of the three given angles, from the interval $0^{0},90^{0}]$.
Only after we had presented our solution [1], it was brought to our attention that this “three-circle problem” was already claimed long ago in [16] by Jakob Steiner, the great geometer of the $19^{\text{th}}$ century – see [3, pp. 468-469, Problem (b)], also [12]. On the bottom of page 162 in [16], Steiner states that he has solved this problem and that the solution will appear in a future monograph. However, it seems almost certain that that monograph was never published.
Notes section of this old Monthly [3, p. 469] attributes the first published solution of this problem to F. Neumann in 1825 [13]. That Monthly note also contains quite a few related sources. For example, the book [11] developes material at great length “to describe a circle which shall cut three given circles at given angles” (pp. 239-241), while [6] uses coaxiability properties of circles to solve “the problem of Steiner,” namely, “to construct a circle meeting three given circles at three given angles.” Furthermore, we were also informed about the master’s degree thesis [14], where a detailed, 30-plus page, solution of the problem of Steiner that utilizes inversion is presented, which most likely was never published.
@@
Our solution of the three-circle problem of Steiner – stated as Problem S in Subsection 3.1 – seems to be different from all that we have been able to find. We use inversive geometry – briefly outlined in Section 4 of this article – with the main tool being constructibility of a circle, the inversion with respect to which transforms two non-intersecting circles into concentric ones – see Subsection 4.2. To solve the problem, we first assume, in Subsection 5.1, that a circle to be constructed exists and perform an inversion  $\mathcal{I}$ of the four circles that transforms two of the three given circles into concentric ones. Then we find a formula for a possible radius of the image – under the inversion $\mathcal{I}$ – of the circle in question and formulas for possible distances from its center to the centers of the images of two of the given circles. Finally, using these three formulas, we construct, in Subsection 5.2, a circle $\mathcal{C}^{\prime}$ that does intersect each of the images of the three given ones at one of the three given angles. As the square of an inversion is the identity map, $\mathcal{I}(\mathcal{C}^{\prime})$ is a circle that intersects each of the three given circles at one of the three given angles, finalizing the solution. Note that the case when each of the three prescribed angles is of $90$ degrees is treated separately.
There is a way to approach the problem of Steiner by using coordinates to obtain a system of equations for the center and radius of the desired circle. But, just doing that and no more does not address the general existence or constructibility of this circle – it simply allows for easy computation and graphing of that circle using software assuming its existence. Some analytic work has been done along these lines – see for example [15]. However, as the authors state in Preface to [7], “The mathematics curriculum in the secondary school normally includes a single one-year course in geometry,” which is usually the student’s “sole exposure to the subject. In contrast, the mathematically minded student has the opportunity of studying elementary algebra, intermediate algebra, and even advanced algebra. It is natural, therefore, to expect a bias in favor of algebra and against geometry. Moreover, misguided enthusiasts lead the student to believe that geometry is “outside the main stream of mathematics” and that analysis or set theory should supersede it.”
Furtermore, the flavor of an analytic treatment would not be consonant with the original role of the problem as an application of the advanced concepts and techniques of synthetic Euclidean geometry that were developed in the late 18th and 19th centuries, such as inversive geometry. We have chosen to keep that spirit alive by presenting a solution to the three-circle problem with prescribed angles that uses these elegant synthetic geometric techniques to show not only that required circle exists, but also how it can be constructed.
Finally, aware of the expository challenge, we set up the explicit goal of producing an account that is self-contained by today’s standards. The older solutions that we have seen are brief and elegant, but are given after lengthy developments of nontrivial synthetic geometric facts and techniques. In order to read them, one must do a lot of backtracking to gather up the necessary pieces before putting them together.
@@ 2.2. Constructibility
First off, we presume that the reader is familiar with the following compass-straightedge constructions:
@@
    Drawing the line through a given point orthogonal to a given line.
@@
    Drawing the line through a given point parallel to a given line.
Let us now try to define the notion of constructibility of an interval:
Definition 2.1.
Let $a_{1},a_{2},\dots,a_{n}$ be drawn intervals. Denote their lengths by the same letters. Let $f$ be a real-valued function of several real variables. An interval of length
$a=f(a_{1},a_{2},\dots,a_{n})$
is called constructible if it can be drawn, using the intervals $a_{1},\dots,a_{n}$, with the help of a straightedge and a compass only. In such a case, the number $a$ is also called constructible.
A comprehensive account on constructibility can be found, for example, in [4]. We shall present just two basic constructions – where we use those in (a) - (c) above without explicit references – that we will need to attain the goal of this article:
Problem 2.1.
Construct the circle that passes through three non-collinear points $A$, $B$, and $C$.
Solution: Utilizing constructions (a) and (b) above, draw perpendicular bisectors of the intervals $[AB]$ and $[BC]$. The point of intersection of these bisectors $O$ is the center of the circle through $A$, $B$, and $C$, and we can draw this circle with a compass since its radius equals $|OA|$.
Problem 2.2.
Given intervals $a$, $b$, and $c$, construct an interval of length $x=\displaystyle\frac{ab}{c}$.
@@
Note that, given an interval $c$ and an angle $\alpha\in(0,\pi/2)$, one can construct a right triangle with the hypotenuse $c$ and an acute angle $\alpha$. Then, as the sides of this triangle are $a=c\sin\alpha$ and $b=c\cos\alpha$, we conclude that the intervals $a$ and $b$ are constructible as soon as $c$ and $\alpha$ are drawn.
3. Statement of the three-circle problem of Steiner
3.1. The Apollonius Three-Circle Problem
Referring to Figure 3, we can state the classical Apollonius three-circle problem as follows:
@@ 3.1. The Apollonius Three-Circle Problem
(Problem A) Given three circles $\mathcal{C}_{1}$, $\mathcal{C}_{2}$, $\mathcal{C}_{3}$, construct a circle $\mathcal{C}$ that is tangent to each of them, that is, that intersects each at one point.
@@
Note that, generically, there are eight different solutions to the Apollonius problem, as the circle $\mathcal{C}$ may touch one, or two, or all three given circles from outside.
Figure 4 below is to state the following generalization of the Apollonius problem, that can be referred to as the three-circle problem of Steiner:
@@
(Problem S) Given three circles  $\mathcal{C}_{1}$, $\mathcal{C}_{2}$, $\mathcal{C}_{3}$ with centers at $O_{1}$, $O_{2}$, $O_{3}$ and radii $r_{1}$, $r_{2}$, $r_{3}$, respectively, construct a circle $\mathcal{C}$ that intersects $\mathcal{C}_{1}$ at an angle $0^{0}\leq\alpha\leq 90^{0}$, $\mathcal{C}_{2}$ at an angle $0^{0}\leq\beta\leq 90^{0}$, and $\mathcal{C}_{3}$ at an angle $0^{0}\leq\gamma\leq 90^{0}$.
@@
4. Inversion and its properties
4.1. Definitions, basic properties and constructions
@@
Points $P\neq O$ and $P^{\prime}\neq O$ are called the inversion images of each other with respect to $\mathcal{C}_{\mathcal{I}}$ if $P$ and $P^{\prime}$ lie on the same ray emanating from $O$ and
@@
This transformation of the plane – minus the point $O$ – is called the inversion with respect to $\mathcal{C}_{\mathcal{I}}$. If we denote this inversion by $\mathcal{I}$, then $\mathcal{I}(P)=P^{\prime}$, $\mathcal{I}(P^{\prime})=P$, hence $\mathcal{I}^{2}(P)=P$ for every $P\neq O$, and $\mathcal{I}(P)=P$ whenever $P\in\mathcal{C}_{\mathcal{I}}$.
Figure 5 gives a way of constructing the point $P^{\prime}$ if $P$ is given and vise versa: If $P$ is given, let $A$ be the intersection of $\mathcal{C}_{\mathcal{I}}$ with the circle for which $[OP]$ is a diameter. Then $PA$ is tangent to $\mathcal{C}_{\mathcal{I}}$, and if $AP^{\prime}\perp OP$, the triangles $AOP$ and $P^{\prime}OA$ are similar, so we have
@@
Next, it can be proven – see [7, Theorem 5.45] – that
@@
    If a circle $\mathcal{C}$ passes through $O$, then $\mathcal{I}(\mathcal{C})$ is a straight line.
Let us now describe a construction that will need to perform repeatedly in the sequel without explicit reference:
Problem 4.1.
Given a circle $\mathcal{C}$ that does not pass through the center of an inversion circle $\mathcal{C}_{\mathcal{I}}$, construct $\mathcal{I}(\mathcal{C})$.
Solution: Pick three points $A,B,C$ on $\mathcal{C}$ and construct the points $\mathcal{I}(A)$, $\mathcal{I}(B)$, $\mathcal{I}(C)$ using the procedure given by Figure 5. Then construct the circle $\mathcal{I}(\mathcal{C})$ as passing through the points $\mathcal{I}(A)$, $\mathcal{I}(B)$, $\mathcal{I}(C)$ as in Problem 2.1.
Our solution of the three-circle problem of Steiner is based on the following facts:
  1. (A)
    An inversion preserves the angles between two circles – see [7, Theorem 5.51]. (Clearly, lines can be included in this statement, as circles with “infinite radii.”)
  2. (B)
    Given two non-intersecting circles $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$, there exists a circle $\mathcal{C}_{\mathcal{I}}$ such that $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ turn into concentric circles under the inversion with respect to $\mathcal{C}_{\mathcal{I}}$ – see [7, Theorem 5.71].
4.2. Constructing a circle $\mathcal{C}_{\mathcal{I}}$ in (B)
Referring to Figure 6, we will describe the process of constructing possible centers for the circle $\mathcal{C}_{\mathcal{I}}$ from the statement (B) above as follows.
@@ 4.2. Constructing a circle $\mathcal{C}_{\mathcal{I}}$ in (B)
Let $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ be non-intersecting circles with centers at $O_{1}$ and $O_{2}$ and radiiradii $r_{1}$ and $r_{2}$, respectively. Constructing a right triangle with the hypotenuse of length $|O_{1}O_{2}|$ and one of the sides of length $|r_{1}-r_{2}|$, we can draw the common outer tangents $A_{1}B_{1}]$ and $A_{2}B_{2}]$ to the circles $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$, respectively. Let $M_{1}$ and $M_{2}$ be the mid-points of segments $A_{1}B_{1}]$ and $A_{2}B_{2}]$, respectively. It is clear that the line $M_{1}M_{2}$ is perpendicular to $O_{1}O_{2}$ and, following the argument in [7, Theorem 2.21], we conclude that $M_{1}M_{2}$ is the radical axis of circles $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$, that is, the tangent intervals to these circles eminating from any point on $M_{1}M_{2}$ are of the same length. Denote by $C$ the point of intersection of the lines $O_{1}O_{2}$ and $M_{1}M_{2}$.
Construct the point $D_{1}\in\mathcal{C}_{1}$ such that $CD_{1}]$ is tangent to $\mathcal{C}_{1}$, as we did in Figure 5 with point $A$. Note that if $D_{2}$ is such a point on $\mathcal{C}_{2}$, then, since the point $C$ lies on the radical axis of $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$, we have $|CD_{2}|=|CD_{1}|$.
Next, draw the circle of radius $|CD_{1}|$ with the center $C$ and denote it by $\mathcal{C}_{0}$. Since $CD_{1}$ and $CD_{2}$ are tangent to $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$, respectively, $\mathcal{C}_{0}$ intersects these circles at $90$-degree angles.
Denote by $\mathcal{O}_{\mathcal{I}_{1}}$ and $\mathcal{O}_{\mathcal{I}_{2}}$ the points of intersection of $\mathcal{C}_{0}$ with $O_{1}O_{2}$, and let $\mathcal{I}$ stand for the inversion with respect to a circle of any non-zero radius centered at, say, $\mathcal{O}_{\mathcal{I}_{2}}$ (which can be clearly replaced by any circle centered at $\mathcal{O}_{\mathcal{I}_{1}}$). By (ii) in Subsection 4.1, $\mathcal{I}(\mathcal{C}_{0})$ is a straight line, which, by (A) above, intersects the circles $\mathcal{I}(\mathcal{C}_{1})$ and $\mathcal{I}(\mathcal{C}_{2})$ at $90$-degree angles, hence passes through each of their centers.
On the other hand, since the center $\mathcal{O}_{\mathcal{I}_{2}}$ lies on the line $O_{1}O_{2}$, this line is invariant with respect to $\mathcal{I}$, and, as $O_{1}O_{2}$ is orthogonal to both $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$, (A) above entails that $O_{1}O_{2}$ is orthogonal to $\mathcal{I}(\mathcal{C}_{1})$ and $\mathcal{I}(\mathcal{C}_{2})$, so, also passes through their centers.
Thus, we have two distinct lines $\mathcal{I}(\mathcal{C}_{0})$ and $O_{1}O_{2}$, each passing through the centers of $\mathcal{I}(\mathcal{C}_{1})$ and $\mathcal{I}(\mathcal{C}_{2})$, implying that these centers coincide, that is, $\mathcal{I}(\mathcal{C}_{1})$ and $\mathcal{I}(\mathcal{C}_{2})$ are concentric.
5. Inversive geometry solution of the Problem of Steiner
An elegant solution of the Apollonius three-circle problem – in a particular case given by Figure 3 – is given via an application of inversion as follows (see, for example [2, pp. 297-299]). First, the radii of the three given circles are enlarged by the same quantity so that two of the resulting circles become tangent at, say, a point $O$. Then an inversion with the center $O$ is performed resulting in these two circles turn into parallel lines. Note that property (B) above grants another striking approach to the Apollonius problem, one that uses a method of Gauss [9].
5.1. Investigation
Assume first that there exists a circle $\mathcal{C}$ that satisfies the conditions in Problem S, as in Figure 4. Let $\mathcal{I}$ be the inversion from 4.2 with the inversion circle $\mathcal{C}_{\mathcal{I}}$ depicted in Figure 7 in green. Apply this inversion to the four other circles in Figure 7, the same as the four circles in Figure 4.